# The ground state energy is not always convex in the number of electrons

### v2

@@ Abstract
We provide the first counter-example showing that the ground state energy of electrons in an external Coulomb potential is not always a convex function of the number of electrons. This property had been conjectured to hold for decades and it plays an important role in quantum chemistry. Our counter-example involves an external potential generated by six nuclei of small fractional charges, placed far away from each other. The ground state energy of 3 electrons is proved to be higher than the average of the energies for 2 and 4 electrons. In addition, we show that the nuclei can bind 2 or 4 electrons, but not 3. This article raises the question of whether the energy convexity really holds for all possible molecules (with nuclei of integer charges).
@@ 1 Introduction
It seems to be an experimental fact that, in an atom or a molecule, the core electrons are more tightly bound to the nuclei than the valence electrons. This property is measured by looking at the ionization energy. Namely, if we remove the electrons of an atom one after another, the corresponding energy cost is seen to grow monotonically, which is a manifestation of an increasing difficulty of pulling out the electrons. In this paper we will provide the very first example of a Coulomb system where the latter property fails. In our example the third electron is more loosely bound to the nuclei than the fourth electron. This certainly goes against our intuition and, in fact, contradicts one of the most famous mathematical conjectures in quantum chemistry.
@@
Figure 1: Ground state energy (eV) of Oxygen as a function of the number $N$ of electrons, according to Parr and Yang9.
@@
It has been known for some time that if Conjecture (4) is valid, this must be due to the particular nature of the Coulomb repulsion between the electrons. In fact, Lieb11 provided a counter-example for $N=2$, with classical hard spheres hopping on 4 sites in space. Lieb’s construction was recently generalized by Ayers17 to the Riesz interaction potential $|\mathbf{r}_{j}-\mathbf{r}_{k}|^{-s}$ with $s>2\log 2/\log 3\approx 1.26$. Hard spheres are recovered in the limit $s\to\infty$. Although $1.26$ is quite close to 1, no counter-example is known at the present for Coulomb. Other negative results concern particles interacting with a repulsive harmonic interaction in a harmonic trap $V$18, 19, the 1D Hubbard model20 or a system in nuclear physics.21
In this article we provide the very first counter-example to Conjecture (4) for the Coulomb interaction. Similarly as for Lieb’s counter-example11, we first consider classical electrons hopping on finitely many sites. Then we extend the counter-example to the quantum case by a perturbation argument.
Our counter-example relies on a recent work22, where we studied the formalism of grand-canonical DFT8 for classical electrons. We found there a situation where the grand-canonical universal density functional does not coincide with its canonical counterpart. In the next section we explain that this is equivalent to the failure of the convexity property (4).
@@ 2 Convexity and grand-canonical DFT
$E_{\rm GC}[V,N]=\!\!\inf_{\begin{subarray}{c}\rho\\ \int_{\mathbb{R}^{3}}\rho=N\end{subarray}}\!\left\{F_{\rm GC}[\rho]+\int_{\mathbb{R}^{3}}\rho(\mathbf{r})V(\mathbf{r})\,{\rm d}\mathbf{r}\right\}$ (13)
@@
$F_{\rm GC}[\rho]=\sup_{V}\left\{E_{\rm GC}[V,N]-\int_{\mathbb{R}^{3}}\rho(\mathbf{r})V(\mathbf{r})\,{\rm d}\mathbf{r}\right\}.$ (14)
@@
The convexity property (4) holds for all $N\in\mathbb{N}$ and all $V$ of the form (3), if and only if we have $F_{\rm GC}[\rho]=F[\rho]$ for all $\rho$ such that $\int_{\mathbb{R}^{3}}\rho(\mathbf{r}){\rm d}\mathbf{r}\in\mathbb{N}$.
Since the validity of the convexity conjecture (4) is equivalent to the equality of two universal functionals in DFT, we can focus on finding a density $\rho$ for which $F[\rho]>F_{\rm GC}[\rho]$. In the low density regime (or equivalently for strong interactions) the two functionals converge to their classical counterparts. It would thus be sufficient to find a counter-example for classical electrons. This is what was accomplished in our recent work22. But let us first quickly mention a positive result, showing that the space dimension plays a crucial role here.
@@ 4 Classical electrons hopping on finitely many sites
Let us now turn to the main result of the paper, that is, the construction of a counter-example to the convexity property (4).
We consider $N$ classical electrons that are only allowed to hop on $K$ distinct sites $\mathbf{R}_{1},...,\mathbf{R}_{K}\in\mathbb{R}^{3}$. We will start by explaining how to formulate the problem for an arbitrary number of sites $K$. Then we will quickly discuss the case of $K=4$ points that was considered by Lieb11 and Ayers17, before we provide the sought-after counter-example for $K=6$.
@@ 4.2 Convexity for $K=4$ sites
Lieb11 found a counter-example to the convexity property for hard-spheres hopping on $K=4$ sites. Ayers17 recently extended this result to the Riesz potential $|\mathbf{r}_{j}-\mathbf{r}_{k}|^{-s}$ for $s>2\log 2/\log 3\approx 1.26$. A natural question is whether one can push $s$ down to 1 and get a counter-example for Coulomb. We can answer the question negatively. The energy is always convex when the electrons are constrained to hop on $K\leqslant 4$ sites!
@@ 4.3 Counter-example for $K=6$ sites
We finally turn to the counter-example. From Theorems 3 and 4, we know that we have to take five points or more in two space dimensions. We will simply place $K=6$ points in a plane, that can be chosen to be the $(x,y)$–plane without loss of generality. We do not consider $K=5$ since we want to work at half-filling and thus need an even number of sites. A counter-example might well exist for $K=5$ too. Following Ref. 22, we take
@@
as displayed in Figure 2. In Ref. 22 we searched for counter-examples using a numerical optimization procedure at half-filling, with random positions for the 6 points. Several of the counter-examples we found were looking a bit like the diamond of Figure 2. To simplify computations, we then restricted the search to this particular shape and this is how we arrived at the mentioned values of $\mathbf{R}^{*}_{1},...,\mathbf{R}^{*}_{6}$. We do not think that there is anything special with this example and many other counter-examples can be found. In fact, the points can be moved a little to suppress the symmetries, but not too much.22
Half-filling means that the density is constrained to be uniform, equal to 1/2 at each site:
$\rho_{\text{h-f}}:=\frac{1}{2}\sum_{k=1}^{6}\delta_{\mathbf{R}_{k}^{*}}$ (23)
@@
We think that the equalities in (25) are an artefact of the particle-hole symmetry. To find a counter-example, we need to slightly go away from half-filling. We thus decided to numerically minimize the “hardness”
@@
Here we have truncated all the values at the fourth decimal. In the following we use these truncated values everywhere, and not the more precise approximation of the minimizer of (26) that we could compute numerically. It is enough to go to the fourth decimal to obtain a counter-example.
@@
5 Counter-example for quantum electrons
To construct a counter-example with quantum electrons, we need first to force them to live close to the points $\mathbf{R}_{j}^{*}$’s in (22) by creating very deep and narrow attractive wells. But then each well has a very negative energy, which we have to compensate by multiplying the Coulomb repulsion by a large constant. After rescaling everything, we see that this amounts to placing the nuclei very far apart.17 We thus consider the following potential
@@ 5 Counter-example for quantum electrons
Theorem 5 (Quantum counter-example)
@@
We recall that the classical energies $E^{\rm cl}[V^{*},N]$ are given in Table 1. Although $E[V^{*}_{\ell},3]$ is lower than in the classical case since the third electron prefers to escape to infinity, we still get the sought-after counter-example
@@
We conclude this section with a remark. In our counter-example the grand-canonical problem for $N=3$ electrons ends up being a strict average of the $2$ and $4$ electron cases. It is not always the case that the grand-canonical $N$-electron problem only involves the $N-1$ and $N+1$ canonical energies. In Theorem 3.16 of Ref. 22 we gave the example of $6^{k}$ points (a kind of multiscale copy of the 6 points in (22) at $k$ different scales), for which the grand-canonical classical problem for $N=6^{k}/2$ is an average of the $N_{-}=(6^{k}-2^{k})/2$ and $N_{+}=(6^{k}+2^{k})/2$ canonical problems. Hence the grand-canonical ground state has a number of electrons that varies a lot around the mean $N$, by $N^{\frac{\log 2}{\log 6}}\gg 1$.
@@ 6 Conclusion
Based on results from our previous work on classical grand-canonical DFT22 and a duality principle, we have provided the very first example of a system of classical electrons for which the ground state energy is not convex in the number of electrons. A perturbation argument then allowed us to construct a quantum counter-example, with 6 nuclei of very small fractional charges placed far apart. It is a very interesting open problem to find a system of real nuclei (that is, of integer charges) for which the convexity property (4) fails. We hope that this paper will stimulate more work in this direction.

### v3

# Ground state energy is not always convex in the number of electrons

@@ Abstract
We provide the first counterexample showing that the ground state energy of electrons in an external Coulomb potential is not always a convex function of the number of electrons. This convexity has been conjectured for decades and plays an important role in quantum chemistry. Our counterexample involves six nuclei with small fractional charges placed far apart. The ground state energy of 3 electrons is shown to be higher than the average of the energies for 2 and 4 electrons. We also show that the nuclei can bind 2 or 4 electrons, but not 3. This article raises the question of whether the energy convexity really holds for all possible molecules (with nuclei of integer charge).
@@ 1 Introduction
It seems to be an experimental fact that in an atom or molecule the core electrons are more tightly bound to the nuclei than the valence electrons. This property is measured by the ionization energy. Namely, if we remove the electrons from an atom one by one, the corresponding energy cost is seen to increase monotonically, which is a manifestation of the increasing difficulty of extracting the electrons. In this paper we will give the first example of a Coulomb system where the latter property fails. In our example, the third electron is more loosely bound to the nuclei than the fourth. This is certainly contrary to our intuition and, in fact, contradicts one of the most famous mathematical conjectures in quantum chemistry.
@@
Figure 1: Experimental ground state energy (eV) of Oxygen as a function of the number $N$ of electrons, with data from the NIST database10. A similar graph can be found in Parr-Yang9.
@@
It has been known for some time that if Conjecture (4) is valid, this must be due to the particular nature of the Coulomb repulsion between the electrons. In fact, Lieb11 provided a counterexample for $N=2$, with classical hard spheres hopping on 4 sites in space. Lieb’s construction was recently generalized by Ayers17 to the Riesz interaction potential $|\mathbf{r}_{j}-\mathbf{r}_{k}|^{-s}$ with $s>2\log 2/\log 3\approx 1.26$. Hard spheres are recovered in the limit $s\to\infty$. Although $1.26$ is quite close to 1, no counterexample is known at the present for Coulomb. Other negative results concern particles interacting with a repulsive harmonic interaction in a harmonic trap $V$18, 19, the 1D Hubbard model20 or a system in nuclear physics.21
In this article we provide the very first counterexample to Conjecture (4) for the Coulomb interaction. Similarly as for Lieb’s counterexample11, we first consider classical electrons hopping on finitely many sites. Then we extend the counterexample to the quantum case by a perturbation argument.
Our counterexample relies on a recent work22, where we studied the formalism of grand-canonical DFT8 for classical electrons. We found there a situation where the grand-canonical universal density functional does not coincide with its canonical counterpart. In the next section we explain that this is equivalent to the failure of the convexity property (4).
@@ 2 Convexity and grand-canonical DFT
$E_{\rm GC}[V,\lambda]=\!\!\inf_{\begin{subarray}{c}\rho\\ \int_{\mathbb{R}^{3}}\rho=\lambda\end{subarray}}\!\left\{F_{\rm GC}[\rho]+\int_{\mathbb{R}^{3}}\rho(\mathbf{r})V(\mathbf{r})\,{\rm d}\mathbf{r}\right\}$ (13)
@@
$F_{\rm GC}[\rho]=\sup_{V}\left\{E_{\rm GC}[V,\lambda]-\int_{\mathbb{R}^{3}}\rho(\mathbf{r})V(\mathbf{r})\,{\rm d}\mathbf{r}\right\}.$ (14)
@@
The convexity property (4) holds for all $N\in\mathbb{N}$ and all $V$ of the form (3), if and only if we have $F_{\rm GC}[\rho]=F[\rho]$ for all $\rho$ such that $\int_{\mathbb{R}^{3}}\rho(\mathbf{r}){\rm d}\mathbf{r}$ is an integer.
Since the validity of the convexity conjecture (4) is equivalent to the equality of two universal functionals in DFT, we can focus on finding a density $\rho$ for which $F[\rho]>F_{\rm GC}[\rho]$. In the low density regime (or equivalently for strong interactions) the two functionals converge to their classical counterparts. It would thus be sufficient to find a counterexample for classical electrons. This is what was accomplished in our recent work22. But let us first quickly mention a positive result, showing that the space dimension plays a crucial role here.
@@ 4 Classical electrons hopping on finitely many sites
Let us now turn to the main result of the paper, that is, the construction of a counterexample to the convexity property (4).
We consider $N$ classical electrons that are only allowed to hop on $K$ distinct sites $\mathbf{R}_{1},...,\mathbf{R}_{K}\in\mathbb{R}^{3}$. We will start by explaining how to formulate the problem for an arbitrary number of sites $K$. Then we will quickly discuss the case of $K=4$ points that was considered by Lieb11 and Ayers17, before we provide the sought-after counterexample for $K=6$.
@@ 4.2 Convexity for $K=4$ sites
Lieb11 found a counterexample to the convexity property for hard-spheres hopping on $K=4$ sites. Ayers17 recently extended this result to the Riesz potential $|\mathbf{r}_{j}-\mathbf{r}_{k}|^{-s}$ for $s>2\log 2/\log 3\approx 1.26$. A natural question is whether one can push $s$ down to 1 and get a counterexample for Coulomb. We can answer the question negatively. The energy is always convex when the electrons are constrained to hop on $K\leqslant 4$ sites!
@@ 4.3 Counterexample for $K=6$ sites
We finally turn to the counterexample. From Theorems 3 and 4, we know that we have to take five points or more in two space dimensions. We will simply place $K=6$ points in a plane, that can be chosen to be the $(x,y)$–plane without loss of generality. We do not consider $K=5$ since we want to work at half-filling and thus need an even number of sites. A counterexample might well exist for $K=5$ too. Following Ref. 22, we take
@@
as displayed in Figure 2. In Ref. 22 we searched for counterexamples using a half-filling numerical optimization procedure with random positions for the 6 points. Several of the counterexamples we found looked a bit like the diamond of Figure 2. To simplify the computations, we then restricted the search to this particular shape and thus arrived at the mentioned values of $\mathbf{R}^{*}_{1},...,\mathbf{R}^{*}_{6}$. We do not think there is anything special about this example, and many other counterexamples can be found. In fact, the points can be moved a little to suppress the symmetries, but not too much.22
Half-filling means that the density is constrained to be uniform, equal to 1/2 at each site,
$\rho_{\text{h-f}}:=\frac{1}{2}\sum_{k=1}^{6}\delta_{\mathbf{R}_{k}^{*}},$ (23)
@@
We think that the equalities in (25) are an artefact of the particle-hole symmetry. To find a counterexample, we need to slightly go away from half-filling. We thus decided to numerically minimize the “hardness”
@@
Here we have truncated all the values at the fourth decimal. In the following we use these truncated values everywhere, and not the more precise approximation of the minimizer of (26) that we could compute numerically. It is enough to go to the fourth decimal to obtain a counterexample.
@@
5 Counterexample for quantum electrons
To construct a counterexample with quantum electrons, we need first to force them to live close to the points $\mathbf{R}_{j}^{*}$’s in (22) by creating very deep and narrow attractive wells. But then each well has a very negative energy, which we have to compensate by multiplying the Coulomb repulsion by a large constant. After rescaling everything, we see that this amounts to placing the nuclei very far apart.17 We thus consider the following potential
@@ 5 Counterexample for quantum electrons
Theorem 5 (Quantum counterexample)
@@
We recall that the classical energies $E^{\rm cl}[V^{*},N]$ are given in Table 1. Although $E[V^{*}_{\ell},3]$ is lower than in the classical case since the third electron prefers to escape to infinity, we still get the sought-after counterexample
@@
We conclude this section with two remarks. First, one can add to the six nuclei any number of even smaller charges placed very far away to obtain a counterexample for $N=3$ and any $K\geqslant 6$. By this procedure, one can also ensure that the total nuclear charge is an integer. Second, in our counterexample the grand-canonical problem for $N=3$ electrons ends up being a strict average of the $2$ and $4$ electron cases. It is also possible to construct a counterexample such that the grand-canonical $N$-electron problem is an average of the $N-J$ and $N+J$ canonical energies with $J$ very large. Indeed, in Theorem 3.16 of Ref. 22 we gave the example of $K=6^{k}$ points (a kind of multiscale copy of the 6 points in (22) at $k$ different scales), for which the grand-canonical classical problem for $N=6^{k}/2$ is an average of the $(6^{k}-2^{k})/2$ and $(6^{k}+2^{k})/2$ canonical problems. Hence the grand-canonical ground state has a number of electrons that varies a lot around the mean $N$, by $N^{\frac{\log 2}{\log 6}}\gg 1$. This gives rise to a quantum counterexample by the same proof as for Theorem 5.
@@ 6 Conclusion
Based on results from our previous work on classical grand-canonical DFT22 and a duality principle, we have provided the very first example of a system of classical electrons for which the ground state energy is not convex in the number of electrons. A perturbation argument then allowed us to construct a quantum counterexample, with 6 nuclei of very small fractional charges placed far apart. It is a very interesting open problem to find a system of real nuclei (that is, of integer charges) for which the convexity property (4) fails. We hope that this paper will stimulate further work in this direction.